Machinery purchased for $70,200 by Sheffield Co. in 2016 was originally estimated to have a life of 8 years with a salvage value of $4,680 at the end of that time. Depreciation has been entered for 5 years on this basis. In 2021, it is determined that the total estimated life should be 10 years with a salvage value of $5,265 at the end of that time. Assume straight-line depreciation.

Prepare the entry to correct the prior years' depreciation, if necessary. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Prepare the entry to record depreciation for 2021. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Machinery related account.)

Account Titles and Explanation

Debit

Credit
Answer #1

a) Journal entry

Account Titles and Explanation

Debit

Credit

No Journal entry 0

b) Journal entry

Accumulated depreciation = (70200-4680/8*5) = 40950

Revised depreciation = (70200-40950-5265/5) = 4797

Account Titles and Explanation

Debit

Credit

Depreciation expense 4797
Accumulated depreciation-Machinery 4797